I need your help on an equation word problem. It is, two angles of a quadrilateral each measure 105 degrees. A third angle measures 5 degrees more than twice the fourth angle. What is,x, the measure of the fourth angle of the quadrilateral? This is what i tried, 2x+2(x+5)=105. x=27.5
If someone can help me as soon as posible please before i go crazy!!!!
Thank you.

Answer by stanbon(75887)

You almost have it.
If the 4th angle is "x", the 3rd angle is 2x+5
EQUATION:
x+2x+5= 360-2(105)
3x+5 = 150
3x = 145
x = 48 1/3 degrees